# In-context Prompt Learning for Test-time Vision Recognition with Frozen Vision-language Model

### v1

@@ Abstract
Existing pre-trained vision-language models, e.g., CLIP, have demonstrated impressive zero-shot generalization capabilities in various downstream tasks. However, the performance of these models will degrade significantly when test inputs present different distributions. To this end, we explore the concept of test-time prompt tuning (TTPT), which enables the adaptation of the CLIP model to novel downstream tasks through only one step of optimization on an unsupervised objective that involves the test sample. Motivated by in-context learning within field of natural language processing (NLP), we propose In-Context Prompt Learning (InCPL) for test-time visual recognition task. InCPL involves associating a new test sample with very few or even just one labeled example as its in-context prompt. As a result, it can reliably estimate a label for the test sample, thereby facilitating the model adaptation process. InCPL first employs a token net to represent language descriptions as visual prompts that the vision encoder of a CLIP model can comprehend. Paired with in-context examples, we further propose a context-aware unsupervised loss to optimize test sample-aware visual prompts. This optimization allows a pre-trained, frozen CLIP model to be adapted to a test sample from any task using its learned adaptive prompt. Our method has demonstrated superior performance and achieved state-of-the-art results across various downstream datasets.
@@ 1 Introduction
Recent advances in vision-language pre-training, such as CLIP [27], have shown a promising direction for developing foundation models for downstream tasks [3]. These foundation models are trained on extensive web-scale data, such as 400 million text-image pairs in the case of CLIP, to align language and vision modalities. When transferring to downstream tasks, the pre-trained model needs to be fine-tuned on a few labeled images for the task domain. Fully fine-tuning a large pre-trained model [3, 19, 6] for each downstream task faces two major challenges in real-world applications. One practical concern is the storage and distribution issue, as maintaining a separate model copy for each task is costly and inflexible, particularly with an expanding number of downstream tasks. Another one is that fully fine-tuning has large probability to destroy initial knowledge provided by the large-scale pre-trained model, increasing the risk of model overfitting.
@@
Unlike the fully fine-tuning to expensively update all model parameters for each downstream task, prompt tuning methods [52, 51] prepend the inputs with learnable parameters that steers the foundation model towards generating the desired outputs. Recent works [52, 51, 2, 15] use training data to tune prompts where the embeddings of prompt are extracted from the model input and are differentiable with respect to the loss function. However, the learned prompts are largely limited to training data, and their performance tend to degrade significantly when the test samples are drawn from a different distribution. Test-time prompt tuning (TTPT) offers a viable solution to the distribution problem through one-step optimization on an unsupervised objective that involves only unlabelled test samples. TTPT flexibly addresses the distribution shift by prompt tuning in a zero-shot manner. For example, TPT [32] exploits the pre-trained vision-language foundation models as better zero-shot learners by maximizing the consistency of predictions across augmented versions of each test sample. However, the learned prompt via existing TTPT methods might be noisy and irrelevant to task-specific distribution. This raises the question: How can we effectively tailor a pre-trained visual-language model (e.g., CLIP) to a new task by judiciously incorporating domain-specific information during testing?
@@
Recent works in natural language processing (NLP) have shown that in-context learning (ICL)  [^1] Without the need to fine-tune any model parameters for downstream tasks, in-context learning involves adding domain-specific input-output pairs, referred to as in-context examples, to a test example. This prompts the model to learn relevant patterns for the test sample, enabling it to perform multiple tasks seamlessly. [5] are effective in adapting foundation models for downstream tasks including knowledge retrieval [38, 41] and multi-choice tasks [24]. Leveraging In-Context Learning (ICL), large language models can tackle novel tasks solely through inference, conditioning on a limited set of input-label pairs to make predictions for new inputs. However, visual model has yet achieve this in-context learning ability. Existing works [36, 20] demonstrates that optimizing an unsupervised test-time objective, involving test samples, can enhance model performance on the target data. Inspired by this, we introduce an innovative approach for vision recognition task during test time without fine-tuning. Our approach prompts the model to acquire domain-specific context information for a test sample using just a few in-context examples.
Figure 1: Illustration of the proposed in-context prompt learning (InCPL) framework. Our approach focuses on identifying unknown objects within a contextual setting using learned prompts, all without the need for extensive fine-tuning of a large pre-trained model. See texts for details.
@@
In this paper, we propose visual In-Context Prompt Learning (InCPL), a method that empowers a pre-trained vision-language model, such as the CLIP model, to leverage in-context examples. Our approach maintains the frozen weights of the vision-language model while enabling gradient back-propagation through it. This allows us to dynamically fine-tune the visual prompt using in-context examples and an unlabeled test sample. To achieve this, we undertake the following steps, as illustrated in Figure 1: (i) We employ a token network to convert textual descriptions into a vision prompt that the vision encoder can comprehend (Figure 1 (a)). (ii) We construct a test sample coupled with in-context candidates to harness the capabilities of the pre-trained large model (Figure 1 (b)). (iii) We design an effective learning strategy to seamlessly integrate the visual prompt with the text prompt (Figure 1 (c)). Through these steps, InCPL enables the CLIP model to adapt to new tasks by leveraging in-context information, all while keeping the model’s core parameters fixed. By harnessing the power of a pre-trained vision-language model, InCPL inherits the robust zero-shot learning capabilities of the CLIP model for downstream tasks on which it was not initially trained. During testing, InCPL seamlessly adapts to new tasks with the assistance of just a few in-context examples and a test sample. Finally, we conduct our experiments on diverse image recognition datasets and evaluate the transfer performance with pre-trained models. Our InCPL outperforms previous methods by a large margin, especially on fine-grained datasets. We also provide comprehensive ablation study and qualitative analysis to understand the effectiveness of InCPL.
@@
Our main contributions can be summarized as follows: 1) We introduce InCPL, a straightforward yet highly effective method for enhancing the CLIP model with in-context examples. This marks the first application of visual in-context prompt learning across a range of downstream tasks. 2) We delve into the effective representation of language descriptions as visual prompt initialization. This involves the utilization of in-context examples and query samples to harness the potential of the vision encoder within the CLIP model. Additionally, we design a learning strategy for visual demonstrations that seamlessly integrate visual prompts with text prompts. 3) Our InCPL approach is rigorously evaluated through extensive experiments, demonstrating superior performance and achieving state-of-the-art (SOTA) results across a diverse set of downstream datasets.
@@ 2 Related work
Prompt Learning in Vision-Language models. Foundational vision-language (V-L) models [27, 14, 49] exploit both visual and textual modalities to encode multi-modal representations. These models are pre-trained a large amount of image-text pairs available online in a self-supervised manner. During pre-training stage, a contrastive loss function is used to pull together the features of paired images and texts while pushing away the unpaired image-text features. V-L models like CLIP [27], FILIP [47] and Florence [48] have demonstrated impressive zero- and few-shot generalization capabilities in a wide range of downstream tasks. Inspired by prompt learning in natural language process (NLP), recent works have proposed to adapt V-L models to downstream task by learning the prompt tokens. CoOp [52] prepends a category name with a prompt “a photo of a” (e.g., “a photo of a cat”), and optimizes continuous set of prompt vectors at its language branch. CoCoOp [51] further makes text prompt conditioned on each input instance (image) and thus enable dynamic text prompt to adapt to each instance. As a visual prompt technique, VP [2, 15] introduces additional random noises as task-specific learnable parameters into input image space. Their prompts are exclusively concentrated within the visual branch. Afterwards, MaPLe [16] extend CoOp [52] and VP [2] to effectively leveraging two kinds of prompts to ensure synergy between vision-language modalities during training stage. However, the learned prompts are limited to the training data distribution and their performances significantly degrade when the model is tested on inputs drawn from a different distribution. To fully delve into the utilization of visual prompts for adapting V-L models, we explore a novel approach based on in-context learning [5] to strongly encourage V-L model adapt to new tasks quickly by using only few in-context examples.
@@
Test-time Adaptation Methods. Test-time adaptation (TTA) [10] focus on challenging scenarios with the presence of only a source model and unlabeled target data are available. This work primarily attempts to address the distribution shift by designing effective test-time objective about test sample. Two similar paradigms are test-time training (TTT)  [36, 20] and domain adaptation (DA) [29, 40]. TTT adds a self-supervised proxy task, such as recognizing rotations of an image, to training objective during training stage and computes an optimization objective at test time, while DA requires the utilization of both source and target data for training with a cross-domain loss function. Without source training data, TTA [31] does not require joint training across losses (TTT) or domain adaptation (DA). The above settings usually needs to fine-tune specific model parameters, which may discard valuable knowledge provided by the pre-trained model and increase the risk of model overfitting. TPT [32] is the first test-time prompt tuning work to enhance the zero-shot generalization of CLIP model by learning adaptive prompts for each test sample at test time. However, TPT focus only on unlabeled test sample, which leads to the prompted (retrieved) knowledge features from CLIP model might deviate from their original meanings within the target distribution.
In-context Learning. In-context learning (ICL) defined by GPT3 [5] is a new paradigm, where autoregressive language model can perform on-the-fly computational reasoning on unseen tasks [23] given prompts and examples serving as context. Flamingo [1] extends the concept of ICL to images and videos by employing language as instruction. [43] conditions the input on a pair of input-output images from the same task to indicate which task to perform. [50] further explore the impact of incontext examples for large vision models. These works  [43, 50] are developed for visual dense prediction but is not applicable to vision and language (V-L) model. Our work follows them but studies visual in-context learning from a different dimension: how to learn prompt for test-time V-L model adaptation with test sample and in-context examples.
@@ 3 Method
3.1 Preliminaries: test-time prompt learning
CLIP Models. In this paper, we focus on generalizing the vision-language pre-trained model (CLIP) [27] to downstream tasks in zero-shot manner while keeping the model parameters frozen. Without supervision from human labels, CLIP can directly learn associated visual-text representations from a large amount of image-text pairs. Specifically, CLIP trains a vision encoder $g_{I}(\cdot)$ and a text encoder $g_{T}(\cdot)$ in contrastive learning to align the embeddings and then match image-text pairs. During testing, zero-shot recognition is accomplished by comparing image features with the class names synthesized by the text encoder.
Test Time Prompt learning. With pre-trained CLIP model, test time prompt learning approaches [32] append learnable prompt tokens at the text encoder during test time, e.g., “a photo of a” is added to every class names to form “a photo of a {class}”. This provides the model very helpful context information about the downstream task. Formally, class label is wrapped within a language prompt, which can be formulated as $\{\bm{t}_{SOS},\bm{P_{\text{{t}}}},\bm{c},\bm{t}_{EOS}\}$. Here, $\bm{c}$ represents the class label, and $\bm{P_{\text{{t}}}}=\{\bm{P_{l}}\}_{l=1}^{4}$ denotes the learnable prompt initialized with word embeddings corresponding to “a photo of a”. $\bm{t}_{SOS}$ and $\bm{t}_{EOS}$ are the learnable start and end token embeddings.
@@ 3.1 Preliminaries: test-time prompt learning
Problem Definition. Given a pre-trained model $f_{\theta_{0}}$ with parameters $\theta_{0}$ and prompt $\bm{P}$, the proposed In-context Test-Time Adaptation (ITTA) aims to adapt $f_{\theta_{0}}$ with only few input-label pairs (known as in-context examples) $(x_{i},y_{i})$ and test samples $x_{t}$ to downstream task. During testing, ITTA fine-tunes a learnable prompt using an optimization objective. This process enhances the model’s generalizability for downstream tasks where test samples lack labels, but in-context examples are accessible. The candidate in-context example set is formed by selecting one image-label pair from each category, followed by random sampling of several pairs from this set to create in-context examples. For a more comprehensive discussion on the selection of examples, we refer to the Section 6.2 in the Appendix. Without requiring source data $(x_{s},y_{s})$, ITTA does not require joint training across domains (domain adaptation) or losses (test time training). In contrast to test-time adaptation, ITTA can provide additional in-context examples as domain-specific context information for target distribution. The difference of ITTA and other adaptations are shown in Table 1.

### v2

@@ Abstract
Current pre-trained vision-language models, such as CLIP, have demonstrated remarkable zero-shot generalization capabilities across various downstream tasks. However, their performance significantly degrades when test inputs exhibit different distributions. In this paper, we explore the concept of test-time prompt tuning (TTPT), which facilitates the adaptation of the CLIP model to novel downstream tasks through a one-step unsupervised optimization that involves only test samples. Inspired by in-context learning in natural language processing (NLP), we propose In-Context Prompt Learning (InCPL) for test-time visual recognition tasks, which empowers a pre-trained vision-language model with labeled examples as context information on downstream task. Specifically, InCPL associates a new test sample with very few labeled examples (sometimes just one) as context information, enabling reliable label estimation for the test sample and facilitating model adaptation. To achieve this, InCPL employs an efficient language-to-vision translator to explore the textual prior information for visual prompt learning. Further, we introduce a context-aware unsupervised loss to optimize visual prompts tailored to test samples. Finally, we design a cyclic learning strategy for visual and textual prompts to ensure mutual synergy across different modalities. This enables a pre-trained, frozen CLIP model to adapt to any task using its learned adaptive prompt. Our method demonstrates superior performance and achieves state-of-the-art results across various downstream datasets.
@@ 1 Introduction
Recent advances in vision-language pre-training, such as CLIP [27], have shown a promising direction for developing foundation models for downstream tasks [2]. These foundation models are trained on extensive web-scale data, such as 400 million text-image pairs in the case of CLIP, to align language and vision modalities. When transferred to downstream tasks, the pre-trained model needs to be fine-tuned on a few labeled images for target domain. Fully fine-tuning a large pre-trained model [2, 19, 5] for each downstream task faces two major challenges in real-world applications. One practical concern is the storage and distribution issue, as maintaining a separate model copy for each task is costly and inflexible, particularly with an expanding number of downstream tasks. Another one is that fully fine-tuning has a large probability of destroying the initial knowledge provided by the large-scale pre-trained model, increasing the risk of model overfitting.
@@
Unlike fully fine-tuning, which expensively updates all model parameters for each downstream task, prompt tuning methods [50, 49] prepend the inputs with learnable parameters that steer the foundation model towards generating the desired outputs. Recent works [50, 49, 1, 15] use training data to tune prompts where the embeddings of prompt are extracted from the model input and are differentiable with respect to the loss function. However, the learned prompts are largely limited to training data, and their performance significantly degrades when the test samples are drawn from a different distribution. Test-time prompt tuning (TTPT) [32, 9] offers a viable solution to the distribution shift problem through one-step optimization on an unsupervised objective that involves only unlabeled test samples. For example, TPT [32] exploits the pre-trained vision-language foundation models as better zero-shot learners by maximizing the consistency of predictions across augmented versions of test samples. However, the learned prompt via existing TTPT methods are noisy and irrelevant to task-specific distribution. This raises the question: How can we effectively adapt a pre-trained visual-language model (e.g., CLIP) to a new task by judiciously incorporating domain-specific information during testing?
@@
Recent works in natural language processing (NLP) have shown that in-context learning (ICL) [^1] Without the need to fine-tune any model parameters for downstream tasks, in-context learning involves adding domain-specific input-output pairs, referred to as in-context examples, to a test example. This prompts the model to learn relevant patterns for the test sample, enabling it to perform multiple tasks seamlessly. [4] is effective in adapting foundation models for downstream tasks including knowledge retrieval [38, 40] and multi-choice tasks [24]. Leveraging in-context learning, large language models (e.g., ChatGPT) can tackle novel tasks solely through inference, conditioning on a limited set of input-label pairs to make predictions for new inputs. However, the visual model has not yet achieved this in-context learning ability. Existing works [36, 20] demonstrate that optimizing an unsupervised test-time objective, involving test samples, can enhance model performance on the target data. Inspired by this, we would like to develop a novel in-context evaluation paradigm that improves the performance of existing prompt learning algorithms in a ChatGPT-like manner. Under this paradigm, the in-context evaluator operates on test samples at test-time with unaltered model parameters, where the model only processes unlabeled test samples and in-context examples.
Figure 1: Illustration of the proposed in-context prompt learning (InCPL) framework. (a) We employ a token network to convert language descriptions into visual prompts, which serve as inputs to the vision encoder of the CLIP model. (b) We construct a test sample coupled with in-context examples and further introduce a context-aware unsupervised loss to optimize visual prompts tailored to test samples. (c) We design a cyclic learning strategy to seamlessly integrate the visual prompt with the text prompt, which effectively retrieves the pre-trained knowledge relevant to test data across different modalities.
@@
Toward this goal, we propose In-Context Prompt Learning (InCPL) that empowers a pre-trained vision-language model, such as the CLIP model, with in-context examples as context information on downstream task. Our approach maintains the frozen weights of the vision-language model while enabling gradient back-propagation through it to learn visual prompt. This enables us to customize visual prompts for different test samples using in-context examples. To achieve this, we undertake the following steps, as illustrated in Figure 1: (i) we employ a token network to convert language descriptions into visual prompts, which serve as inputs to the vision encoder of the CLIP model (Figure 1 (a)). (ii) We construct a test sample coupled with in-context examples and further introduce a context-aware unsupervised loss to optimize visual prompts tailored to test samples (Figure 1 (b)). (iii) We design a cyclic learning strategy to seamlessly integrate the visual prompt with the text prompt, which effectively retrieves the pre-trained knowledge relevant to test data across different modalities (Figure 1 (c)). Through these steps, InCPL enables the CLIP model to adapt to new tasks by leveraging domain-specific information, while keeping the model parameters frozen. By harnessing the power of a pre-trained vision-language model, our InCPL inherits the robust zero-shot learning capabilities of the CLIP model for downstream tasks on which it was not initially trained. As a result, our approach can effectively adapt pre-trained model to new tasks with learned visual prompts. Finally, we conduct our experiments on diverse image recognition datasets and evaluate the transfer performance with pre-trained models. Experimental results show that our InCPL outperforms previous methods by large margins, especially on fine-grained datasets. We also provide comprehensive ablation study and qualitative analysis to understand the effectiveness of InCPL.
@@
Our main contributions can be summarized as follows: 1) We propose a straightforward yet highly in-context learning framework to instruct the CLIP model with in-context examples as context information, which enhances the performance of prompt learning algorithms in a ChatGPT-like manner. 2) We propose an efficient language-to-vision translator to leverage the textual prior information from the language modality, aiding in the learning of visual prompts. 3) We design a cyclic learning strategy for visual and textual prompts to ensure mutual synergy across different modalities. 4) Our InCPL approach is rigorously evaluated through extensive experiments, demonstrating superior performance and achieving state-of-the-art (SOTA) results across a diverse set of downstream datasets.
@@ 2 Related work
Prompt Learning in Vision-language Models. Foundational vision-language (V-L) models [27, 14, 47] exploit both visual and textual modalities to encode multi-modal representations. These models are pre-trained on a large amount of image-text pairs available online in a self-supervised manner. During the pre-training stage, a contrastive loss function is used to pull together the features of paired images and texts while pushing away unpaired image-text features. V-L models like CLIP [27], FILIP [45] and Florence [46] have demonstrated impressive zero- and few-shot generalization capabilities in a wide range of downstream tasks. Inspired by prompt learning in natural language processing (NLP), recent works have proposed to adapt V-L models to downstream tasks by learning the prompt tokens. CoOp [50] prepends a category name with the prompt “a photo of a” (e.g., “a photo of a cat”), and optimizes continuous set of prompt vectors at its language branch. CoCoOp [49] further makes textual prompt conditioned on each input instance (image) and thus enables dynamic text prompt to adapt to each instance. As a visual prompt technique, VP [1, 15] introduces additional random noise as task-specific learnable parameters into input image space. Their prompts are exclusively concentrated within the visual branch. Afterwards, MaPLe [16] extends CoOp [50] to effectively leveraging two kinds of prompts to ensure synergy between vision-language modalities during training stage. However, these learned prompts are limited to the training data distribution and their performances significantly degrade when the model is tested on inputs drawn from a different distribution. To fully delve into the utilization of visual prompts for adapting V-L models, we explore a novel approach based on in-context learning [4] to strongly encourage V-L model adapt to new tasks quickly by using only few in-context examples.
@@
Test-time Adaptation Methods. Test-time adaptation (TTA) [10] focus on challenging scenarios where only a source model and unlabeled target data are available. This work primarily attempts to address the distribution shift by designing effective test-time objective for test sample. Two similar paradigms are test-time training (TTT)  [36, 20] and domain adaptation (DA) [29, wu2019joint]. TTT adds a self-supervised proxy task, such as recognizing rotations of an image, to training objective during training stage and computes an optimization objective at test time, while DA requires the utilization of both source and target data for training with a cross-domain loss function. Without source training data, TTA [31] does not require joint training across losses (TTT) or domain adaptation (DA). The above settings usually need to fine-tune specific model parameters, which may discard valuable knowledge provided by the pre-trained model and increase the risk of model overfitting. TPT [32] is the first test-time prompt tuning work to enhance the zero-shot generalization of CLIP model by learning adaptive prompts for each test sample at test time. DiffTPT [9] extend test-time prompt tuning by leveraging pre-trained diffusion models to augment the diversity of test samples. However, TPT [32] and DiffTPT [9] focus only on unlabeled test sample, which leads to the prompted (retrieved) knowledge features from CLIP model deviates from their original meanings within the target distribution.
In-context Learning. In-context learning (ICL) defined by GPT3 [4] is a new paradigm, where autoregressive language model can perform on-the-fly computational reasoning on unseen tasks [23], given prompts and examples serving as context. Following this idea, [43] and [42] use task-specific examples to help model understand the underlying task and perform the same task on a new input. [43] enable in-context learning in diffusion-based generative models, while [42] allows the model to rapidly adapt to visual dense prediction task with labeled examples. However, these works only employ image-label pairs as task prompts and train overall model on a wide range of tasks. Unlike existing methods, our approach maintains the pre-trained model parameters and tunes the input prompts to retrieve prior knowledge relevant to the test samples. Consequently, our method is more efficient.
2.1 Preliminaries and Problem Definition
CLIP Models. In this paper, we focus on generalizing the pre-trained vision-language model (i.e., CLIP) [27] to downstream tasks in zero-shot manner while keeping the model parameters frozen. Without supervision from human labels, CLIP can directly learn associated visual-text representations from a large amount of image-text pairs. Specifically, CLIP trains a vision encoder $g_{I}(\cdot)$ and a text encoder $g_{T}(\cdot)$ in contrastive learning to align the embeddings and then match image-text pairs. During testing stage, zero-shot recognition is accomplished by comparing image features with the class features synthesized by the text encoder on class names.
Test-time Prompt Learning. With pre-trained CLIP model, test-time prompt learning approaches [32] append learnable prompt tokens at the text encoder during test time, i.e., “a photo of a” is added to every class name to form “a photo of a {class}”. This provides the model very helpful context information about the downstream task. Formally, class label is wrapped within a language prompt, which can be formulated as $\{\bm{t}_{SOS},\bm{P_{\text{{t}}}},\bm{c},\bm{t}_{EOS}\}$. Here, $\bm{c}$ represents the class label, and $\bm{P_{\text{{t}}}}=\{\bm{P_{l}}\}_{l=1}^{4}$ denotes the learnable prompt initialized with word embeddings‘ corresponding to “a photo of a”. $\bm{t}_{SOS}$ and $\bm{t}_{EOS}$ are the learnable start and end token embeddings.
@@ 3 Method
Problem Definition. Given a pre-trained model $f_{\theta_{0}}$ with parameters $\theta_{0}$ and prompt $\bm{P}$, the proposed In-context Test-Time Adaptation (ITTA) aims to adapt $f_{\theta_{0}}$ to downstream task with only few input-label pairs (known as in-context examples) $(x_{i},y_{i})$ and test samples $x_{t}$. During testing stage, ITTA fine-tunes a learnable prompt using an unsupervised optimization objective where test samples lack labels, but in-context examples are accessible. The candidate in-context example set is formed by selecting one image-label pair from each category, followed by random sampling of several pairs from this set to create in-context examples. For a more comprehensive discussion on the selection of examples, we refer to the Section 4.5. In the absence of source domain data $(x_{s},y_{s})$, domain adaptation and test-time training, which require joint training across different domains, are not feasible. While test-time adaptation also does not rely on source domain data, our ITTA can provide additional labeled examples as context information for the test domain. The detailed differences between ITTA and other adaptation methods are shown in Table 1.
3.1 Visual Prompt Input with Images
Figure 2: The comparison between our language-aware visual prompt approach and the patched, padded, and token-based visual prompt methods.
The visual prompt approach enables the learning of image perturbations (such as noisy boxes or rectangles) for a frozen model. This enables the model, when prompted with these perturbations, to perform a new task. As depicted in Figure 2 (a) and (b), recent studies [1, 15] have introduced task-specific learnable parameters as visual prompts. These prompts are prepended into the input image space and are learned during the fine-tuning stage. However, these parameters are typically initialized with random noise. We argue that a general text description (e.g., a photo of a) provides richer visual context information compared to random noise. This is because linguistic words can be perceived as interpretable lexical tokens that can be easily translated into visual content comprehensible to the vision encoder. This is especially advantageous during the test-time adaptation, which lacks effective supervision information.
A straightforward approach to implementing a visual prompt is to combine patch image embedding with learnable parameter vectors initialized with word embeddings. However, this design not only lacks a seamless connection between the vision and language modalities but also faces the challenge of differences in dimensionality between these two types of embeddings. We propose an efficient language-to-vision translation design to leverage the textual prior information from the language modality, aiding in the learning of visual prompts. Specifically, we introduce a lightweight neural network, known as Token-Net, to convert textual descriptions into a vision prompt that the vision encoder can comprehend. Token-Net is responsible for generating a conditional token vector for each input image, which is then integrated with the context visual vectors. For a illustration of this architecture, please refer to Figure 3.